OpenStudy (anonymous):

Connie is factoring 0.36t^28 – 25 by using the rule a^2 – b^2 = (a + b)(a – b). What will she use for the value of a? (Points : 1) 0.6 0.36 0.6t^14 0.06t^14

OpenStudy (whpalmer4):

If \[a^2 = 0.36t^{28}\]then \[a=\]

OpenStudy (anonymous):

0.36 * 28

OpenStudy (whpalmer4):

take the square root of both sides: \[\sqrt{a^2} = \sqrt{0.36t^{28}}\]\[a=\sqrt{0.36*t^{28}}\] \[a=\sqrt{0.36}*\sqrt{t^{28}}\]do you know what \(\sqrt{0.36}=\)

OpenStudy (anonymous):

no....... :(

OpenStudy (whpalmer4):

it is going to be a number, let's call it \(x\), where \(x*x = 0.36\) Is \(x\) going to be larger or smaller than 1?

OpenStudy (anonymous):

SMALLER

OpenStudy (anonymous):

o.6

OpenStudy (whpalmer4):

Right, \[0.6*0.6 = 0.36\] Now what about the \[\sqrt{t^{28}}\]what does that equal?

OpenStudy (whpalmer4):

sorry, got tied up on a phone call